# WizardMath: Empowering Mathematical Reasoning for Large Language Models via Reinforced Evol-Instruct

### v2

@@ 1 Introduction
Figure 1: A diagram illustrating the three steps of our Reinforcement Learning from Evol-Instruct Feedback (RLEIF)
Chain-of-thought (CoT) (Wei et al., 2022) proposes to design better prompts to generate step-by-step solutions, which can lead to improved performance. Self-Consistency (Wang et al., 2022) also achieves remarkable performance on many reasoning benchmarks, which generates several possible answers from the model and selects the correct one based on majority vote (Fu et al., 2022). Llemma (Azerbayev et al., 2023) and MathPile (Wang et al., 2023f) continue pretraining LLMs with math corpus to improve domain capacity. MetaMath (Yu et al., 2023b) and Xwin-Math (Li et al., 2024a) bootstraps mathematical questions by augmenting the question from multiple perspectives. MAmmoTH  (Yue et al., 2023) and TORA (Gou et al., 2023) presents a unique hybrid of CoT and program-of-thought (PoT) to ensure extensive coverage of diverse fields in math. Recently, Evol-Instruct is an effective method for large-scale data synthesis using LLMs. It has been widely verified and proven to be effective in enhancing the model’s instruction following capability. It employs In-depth Evolving and In-breadth Evolving to automate the generation of diverse and complex open-domain instructions using LLMs, instead of relying on human-crafted instruction datasets. In-depth Evolving incrementally enhances instruction complexity by introducing additional constraints, deepening, concretizing, increasing reasoning steps, and complicating input. In-breadth Evolving focuses on improving topic diversity and dataset richness by creating entirely new instructions. To enhance the correctness of each step in the model’s generation process, (Wang et al., 2024a; Chen et al., 2024a; Lightman et al., 2023) finds that process supervision with reinforcement learning significantly outperforms outcome supervision for solving challenging MATH problems.
@@
We perform experiments on two widely used mathematical reasoning benchmarks, namely GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) covering math problems from grade to high school levels, the results show that our WizardMath outperforms all other open-source LLMs at the same model size, achieving state-of-the-art performance. For instance, WizardMath-70B significantly outperforms MetaMath-70B by a significant margin on GSM8k (92.8 vs. 82.3) and on MATH (58.6 vs. 26.6). Specifically, WizardMath-Mistral-7B observed a substantial improvement in pass@1 with an increase of +12.8 (90.7. vs. 77.9) on GSM8k, and +26.8 (55.4 vs. 28.6) on MATH compared to MetaMath-Mistral-7B. Notably, our 70B model even also significantly surpasses those powerful proprietary LLMs, such as GPT-3.5-Turbo, Claude 2 (Bai et al., 2022), Mistral Medium (Jiang et al., 2024a), Gemini-Pro  (Team, 2023), PaLM-2 (Anil et al., 2023) and GPT-4-early-version.
@@ 2 Related Work
Large Language Models For Mathematical reasoning. NLP models face challenges with complex reasoning, including mathematical (Long et al., 2024; Zhang et al., 2024c; Xia et al., 2024), common-sense (Talmor et al., 2019). Significant research focuses on Mathematical Word Problems (MWP), which demand understanding of mathematical concepts and multi-step reasoning (Zheng et al., 2023; Zhao et al., 2023; Yuan et al., 2023a). Models are tested on various MWP benchmarks  (Roy & Roth, 2015; Hendrycks et al., 2021). Techniques like Chain-of-Thought Prompting (Wei et al., 2022), Least-to-Most prompting (Zhou et al., 2022), and Complex CoT (Fu et al., 2022) enhance reasoning by introducing multiple steps and breaking problems into sub-problems. There are some models aimed at improving math CoT reasoning skills such as MetaMath (Yu et al., 2023b), MathScale (Tang et al., 2024), Xwin-Math (Li et al., 2024a), DART-Math (Tong et al., 2024) etc. Some models enhance mathematical reasoning by integrating python tools, such as TORA (Gou et al., 2023), MAmmoTH (Yue et al., 2023), Openmathinstruct (Toshniwal et al., 2024), NuminaMath (Li et al., 2024c) etc. In our work, we mainly improve the CoT reasoning ability of mathematics without using external Python tools.
@@
Reinforcement Learning for Large Language Models. State-of-the-art models often display logical errors and illusions, particularly in domains requiring complex, multi-step reasoning, leading to significant challenges (Bubeck et al., 2023; Maynez et al., 2020). Strategies such as training reward models help discriminate between desirable and undesirable outputs (Lightman et al., 2023; Wu et al., 2023b; Chen et al., 2024b). Historically, outcome-based approaches focused on algorithmic tasks (Li et al., 2016; Cai et al., 2017; Yu et al., 2023a), while recent research demonstrates the efficacy of reward models or validators in enhancing model performance (Cobbe et al., 2021; Wang et al., 2023c; d; Li et al., 2022a). Reward models have also been incorporated into reinforcement learning pipelines and employed in rejection sampling to align Large Language Models (LLMs) with human preferences (Shen et al., 2021; Bai et al., 2022; Yuan et al., 2023c; Dong et al., 2023; Song et al., 2023; Touvron et al., 2023b; Rafailov et al., 2024; Meng et al., 2024). A contrast is drawn between outcome-supervised and process-supervised reward models, with the latter being more effective at addressing discrepancies arising from incorrect reasoning paths leading to correct outcomes (Uesato et al., 2022; Zelikman et al., 2022; Creswell et al., 2022). Recent advances have promoted process-based supervision through manual annotation, significantly benefiting LLMs over outcome-based approaches (Lightman et al., 2023; Wang et al., 2024a; Sun et al., 2024; Chen et al., 2024a; Wang et al., 2024b; Zhang et al., 2024a). In this paper, we leverage AI models like ChatGPT to automatically offer process annotation to improve the efficiency of this research line.
@@ 4.1 Experimental Setup
KPMath-Plus (Huang et al., 2024b) Mistral-v0.1 7B 82.1 46.8
@@
KPMath-Plus (Huang et al., 2024b) DeepSeekMath 7B 83.9 48.8
@@
KPMath-Plus (Huang et al., 2024b) Llama 2 13B 81.6 41.0
@@
KPMath-Plus (Huang et al., 2024b) Llama-2 70B 87.4 48.6
@@ 4.3 ANALYSIS
ORM v.s. PRM; Human v.s. AI. The Table 5 presents the performance of different answer reward methods for LLMs in terms of pass@1. As is shown: 1) Our step-by-step PRM significantly enhances the performance of both Llama and Mistral based SFT models. Specifically, the Mistral-7B powered by our PRM achieves 87.2% and 52.7% on GSM8k and MATH respectively. 2) PRM models consistently outperforms ORM on both GSM8k and MATH, indicating the effectiveness of step-by-step supervision. 3) The PRM trained on our fully AI-labeled data outperforms both the manually annotated PRM800k and Math-Shepherd, which utilizes MCTS tree search for annotation. When training WizardMath-Mistral-SFT with PPO, our PRM improves upon PRM800k by 1.8% and Math-Shepherd by 1.1% on GSM8k, while surpassing PRM800k by 1.9% and Math-Shepherd by 2.4% on MATH. This demonstrates powerful AI can also provide good process supervision quality, highlighting the effectiveness of utilizing AI to construct PRM training data.
@@ A.5 Similarity Checking And Data Filtering
A.6 Detail Works
Large Language Models. LLMs have significantly advanced Natural Language Processing, with models like OpenAI’s GPT Series  (Brown et al., 2020a; OpenAI, 2023), Anthropic’s Claude (Bai et al., 2022), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Gemini (Team, 2023), and Gemma (Team et al., 2024) featuring billions of parameters and trained on massive textual datasets. The AI field has also seen a rise in open-source LLMs such as Mistral (Jiang et al., 2023), Llama Series  (Touvron et al., 2023a; b; Dubey et al., 2024; Taylor et al., 2022), DeepSeek (Bi et al., 2024; Shao et al., 2024), Qwen (Bai et al., 2023; Yang et al., 2024) etc. (Zeng et al., 2022; Penedo et al., 2023; Scao et al., 2022). Notably, Llama serves as a foundational model for supervised fine-tuning, leading to the development of models like Alpaca, Vicuna, Guanaco, and Orca (Taori et al., 2023; Chiang et al., 2023; Dettmers et al., 2023; Mukherjee et al., 2023).
Large Language Models For Mathematical reasoning. NLP models face challenges with complex reasoning, including mathematical (Lu et al., 2022; Frieder et al., 2023; Long et al., 2024; Zhang et al., 2024c; Xia et al., 2024), common-sense (Talmor et al., 2019; Geva et al., 2021). Significant research focuses on Mathematical Word Problems (MWP), which demand understanding of mathematical concepts and multi-step reasoning (Koncel-Kedziorski et al., 2016; Patel et al., 2021; Lan et al., 2022; Cobbe et al., 2021; Jie et al., 2022; Yuan et al., 2023b; Fu et al., 2023b; Zheng et al., 2023; Zhao et al., 2023; Wang et al., 2023b; Imani et al., 2023; Yuan et al., 2023a; Wang et al., 2023e; Ahn et al., 2024). Models are tested on various MWP benchmarks  (Hosseini et al., 2014; Roy & Roth, 2015; Koncel-Kedziorski et al., 2015; Patel et al., 2021; Cobbe et al., 2021; Ling et al., 2017; Hendrycks et al., 2021; Xia et al., 2024; Huang et al., 2024a; Zhang et al., 2024c; Li et al., 2024d; Anand et al., 2024). Techniques like Chain-of-Thought Prompting (Wei et al., 2022), Least-to-Most prompting (Zhou et al., 2022), and Complex CoT (Fu et al., 2022) enhance reasoning by introducing multiple steps and breaking problems into sub-problems. There are some models aimed at improving math CoT reasoning skills such as MetaMath (Yu et al., 2023b), MathScale (Tang et al., 2024), Xwin-math (Li et al., 2024a) etc. (Liu et al., 2023; Liu & Yao, 2024; Jiang et al., 2024b; Xiong et al., 2024; Chan et al., 2024; Lin et al., 2024; Huang et al., 2024b; Mitra et al., 2024b; Yuan et al., 2023a; Wu et al., 2024; Mitra et al., 2024a; Zhang et al., 2024e; Fu et al., 2023a; Ni et al., 2024; Tian et al., 2024; Zhang et al., 2024b; Zeng et al., 2024; Shao et al., 2024; Feng et al., 2023; Wu et al., 2023a; Tong et al., 2024; Yang et al., 2024; Zhou et al., 2024; Chen et al., 2024c; Zhang et al., 2024d; Su et al., 2024). Some models enhance mathematical reasoning by integrating python tools, such as TORA (Gou et al., 2023), MAmmoTH (Yue et al., 2023), Openmathinstruct (Toshniwal et al., 2024), etc. (Lu et al., 2024; Li et al., 2024c; Yu et al., 2024; Wang et al., 2023a; Li et al., 2024b)
Large Language Models For Reinforcement Learning. State-of-the-art models often display logical errors and illusions, particularly in domains requiring complex, multi-step reasoning, leading to significant challenges (Bubeck et al., 2023; Maynez et al., 2020). Strategies such as training reward models help discriminate between desirable and undesirable outputs (Lightman et al., 2023; Wu et al., 2023b; Chen et al., 2024b). Historically, outcome-based approaches focused on algorithmic tasks (Graves et al., 2014; Reed & De Freitas, 2015; Li et al., 2016; Cai et al., 2017; Yu et al., 2023a), while recent research demonstrates the efficacy of reward models or validators in enhancing model performance (Cobbe et al., 2021; Wang et al., 2023c; d; Li et al., 2022a; b). Reward models have also been incorporated into reinforcement learning pipelines and employed in rejection sampling to align Large Language Models (LLMs) with human preferences (Ziegler et al., 2019; Stiennon et al., 2020; Nakano et al., 2021; Ouyang et al., 2022; Nichols et al., 2020; Shen et al., 2021; Bai et al., 2022; Yuan et al., 2023c; Dong et al., 2023; Song et al., 2023; Touvron et al., 2023b). A contrast is drawn between outcome-supervised and process-supervised reward models, with the latter being more effective at addressing discrepancies arising from incorrect reasoning paths leading to correct outcomes (Uesato et al., 2022; Zelikman et al., 2022; Creswell et al., 2022). Recent advances have promoted process-based supervision through manual annotation, significantly benefiting LLMs over outcome-based approaches (Lightman et al., 2023; Zhu et al., 2023; Ni et al., 2022; Wang et al., 2024a; Sun et al., 2024; Chen et al., 2024a; Wang et al., 2024b; Zhang et al., 2024a; b). In this paper, we leverage AI models like ChatGPT to automatically offer annotation to improve the efficiency of this research line.
@@ A.6 Detail Works
KPMath-Plus (Huang et al., 2024b) Mistral-v0.1 7B 82.1 46.8
@@
KPMath-Plus (Huang et al., 2024b) DeepSeekMath 7B 83.9 48.8
@@
KPMath-Plus (Huang et al., 2024b) Llama 2 13B 81.6 41.0
@@
KPMath-Plus (Huang et al., 2024b) Llama-2 70B 87.4 48.6
@@
A.7 compare our approach with the advanced method for SFT using synthesized data, such as DartMath
As the volume of training data increases, WizardMath-Evol-Instruct consistently improves its performance on the GSM8k and MATH benchmarks, exhibiting a slightly higher growth rate than DART-Math. In the initial stages, WizardMath slightly underperforms compared to DART-Math. This advantage may stem from DART-Math being distilled from DeepSeekMath-RL, an advanced mathematical reasoning model pre-trained on 120B high-quality mathematical tokens, showcasing exceptional proficiency in mathematical reasoning. However, once the dataset exceeds 60k, its performance begins to surpasse DART-Math. At a data scale of 390k, WizardMath slightly outperforms DART-Math by 2%–3% on GSM8k and by 5%–6% on MATH. Additionally, WizardMath-Evol-Instruct consistently exceeds MetaMath at the same data scales, achieving increases of 3%–6% on GSM8k and 15%–20% on MATH. This performance gain is attributed to the efficiency of Math Evol-Instruct’s upward and downward evolution processes. These findings demonstrate that our Math Evol-Instruct method is also as scalable and effective as DART-Math for the large-scale synthetic data.

### v3

@@ 1 Introduction
Figure 1: A diagram illustrating the three steps of our Reinforcement Learning from Evol-Instruct Feedback (RLEIF). For a detailed explanation of the training pipeline, refer to Appendix  A.6
Chain-of-thought (CoT) (Wei et al., 2022) proposes to design better prompts to generate step-by-step solutions, which can lead to improved performance. Self-Consistency (Wang et al., 2022) also achieves remarkable performance on many reasoning benchmarks, which generates several possible answers from the model and selects the correct one based on majority vote (Fu et al., 2022). Llemma (Azerbayev et al., 2023) and MathPile (Wang et al., 2023c) continue pretraining LLMs with math corpus to improve domain capacity. MetaMath (Yu et al., 2023b) and Xwin-Math (Li et al., 2024a) bootstraps mathematical questions by augmenting the question from multiple perspectives. MAmmoTH  (Yue et al., 2023) and TORA (Gou et al., 2023) presents a unique hybrid of CoT and program-of-thought (PoT) to ensure extensive coverage of diverse fields in math. Recently, Evol-Instruct is an effective method for large-scale data synthesis using LLMs. It has been widely verified and proven to be effective in enhancing the model’s instruction following capability. It employs In-depth Evolving and In-breadth Evolving to automate the generation of diverse and complex open-domain instructions using LLMs, instead of relying on human-crafted instruction datasets. In-depth Evolving incrementally enhances instruction complexity by introducing additional constraints, deepening, concretizing, increasing reasoning steps, and complicating input. In-breadth Evolving focuses on improving topic diversity and dataset richness by creating entirely new instructions. To enhance the correctness of each step in the model’s generation process, (Wang et al., 2024a; Chen et al., 2024a; Lightman et al., 2023) finds that process supervision with reinforcement learning significantly outperforms outcome supervision for solving challenging MATH problems.
@@
We perform experiments on two widely used mathematical reasoning benchmarks, namely GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) covering math problems from grade to high school levels, the results show that our WizardMath outperforms all other open-source LLMs at the same model size, achieving state-of-the-art performance. For instance, WizardMath-70B significantly outperforms MetaMath-70B by a significant margin on GSM8k (92.8 vs. 82.3) and on MATH (58.6 vs. 26.6). Specifically, WizardMath-Mistral-7B observed a substantial improvement in pass@1 with an increase of +12.8 (90.7. vs. 77.9) on GSM8k, and +26.8 (55.4 vs. 28.6) on MATH compared to MetaMath-Mistral-7B. Notably, our 70B model even also significantly surpasses those powerful proprietary LLMs, such as GPT-3.5-Turbo, Claude 2 (Bai et al., 2022), Mistral Medium (Jiang et al., 2024), Gemini-Pro  (Team, 2023), PaLM-2 (Anil et al., 2023) and GPT-4-early-version.
@@ 2 Related Work
Large Language Models For Mathematical reasoning. NLP models face challenges with complex reasoning, including mathematical (Long et al., 2024; Zhang et al., 2024b; Xia et al., 2024), common-sense (Talmor et al., 2019). Significant research focuses on Mathematical Word Problems (MWP), which demand understanding of mathematical concepts and multi-step reasoning (Zheng et al., 2023; Zhao et al., 2023; Yuan et al., 2023a). Models are tested on various MWP benchmarks  (Roy & Roth, 2015; Hendrycks et al., 2021). Techniques like Chain-of-Thought Prompting (Wei et al., 2022), Least-to-Most prompting (Zhou et al., 2022), and Complex CoT (Fu et al., 2022) enhance reasoning by introducing multiple steps and breaking problems into sub-problems. There are some models aimed at improving math CoT reasoning skills such as MetaMath (Yu et al., 2023b), MathScale (Tang et al., 2024), Xwin-Math (Li et al., 2024a), DART-Math (Tong et al., 2024) etc. Some models enhance mathematical reasoning by integrating python tools, such as TORA (Gou et al., 2023), MAmmoTH (Yue et al., 2023), Openmathinstruct (Toshniwal et al., 2024), NuminaMath (Li et al., 2024b) etc. In our work, we mainly improve the CoT reasoning ability of mathematics without using external Python tools.
@@
Reinforcement Learning for Large Language Models. State-of-the-art models often display logical errors and illusions, particularly in domains requiring complex, multi-step reasoning, leading to significant challenges (Bubeck et al., 2023; Maynez et al., 2020). Strategies such as training reward models help discriminate between desirable and undesirable outputs (Lightman et al., 2023; Wu et al., 2023; Chen et al., 2024b). Historically, outcome-based approaches focused on algorithmic tasks (Li et al., 2016; Cai et al., 2017; Yu et al., 2023a), while recent research demonstrates the efficacy of reward models or validators in enhancing model performance (Cobbe et al., 2021; Wang et al., 2023a; b; Li et al., 2022). Reward models have also been incorporated into reinforcement learning pipelines and employed in rejection sampling to align Large Language Models (LLMs) with human preferences (Shen et al., 2021; Bai et al., 2022; Yuan et al., 2023b; Dong et al., 2023; Song et al., 2023; Touvron et al., 2023b; Rafailov et al., 2024; Meng et al., 2024). A contrast is drawn between outcome-supervised and process-supervised reward models, with the latter being more effective at addressing discrepancies arising from incorrect reasoning paths leading to correct outcomes (Uesato et al., 2022; Zelikman et al., 2022; Creswell et al., 2022). Recent advances have promoted process-based supervision through manual annotation, significantly benefiting LLMs over outcome-based approaches (Lightman et al., 2023; Wang et al., 2024a; Sun et al., 2024; Chen et al., 2024a; Wang et al., 2024b; Zhang et al., 2024a). In our study, we leverage AI models like ChatGPT to automatically offer process annotation to improve the efficiency of this research line.
@@ 4.1 Experimental Setup
KPMath-Plus (Huang et al., 2024) Mistral-v0.1 7B 82.1 46.8
@@
KPMath-Plus (Huang et al., 2024) DeepSeekMath 7B 83.9 48.8
@@
KPMath-Plus (Huang et al., 2024) Llama 2 13B 81.6 41.0
@@
KPMath-Plus (Huang et al., 2024) Llama-2 70B 87.4 48.6
@@ 4.3 ANALYSIS
ORM v.s. PRM; Human v.s. AI. Table 5 presents the performance of different answer reward methods for LLMs in terms of pass@1. As is shown: 1) Our step-by-step PRM significantly enhances the performance of both Llama and Mistral based SFT models. Specifically, the Mistral-7B powered by our PRM achieves 87.2% and 52.7% on GSM8k and MATH respectively. 2) PRM models consistently outperforms ORM on both GSM8k and MATH, indicating the effectiveness of step-by-step supervision. 3) The PRM trained on our fully AI-labeled data outperforms both the manually annotated PRM800k and Math-Shepherd, which utilizes MCTS tree search for annotation. When training WizardMath-Mistral-SFT with PPO, our PRM improves upon PRM800k by 1.8% and Math-Shepherd by 1.1% on GSM8k, while surpassing PRM800k by 1.9% and Math-Shepherd by 2.4% on MATH. This demonstrates powerful AI can also provide good process supervision quality, highlighting the effectiveness of utilizing AI to construct PRM training data.
@@ A.5 Similarity Checking And Data Filtering
A.6 Detailed explanation of our method flow.
We offer a detailed clarification of our method flow , including the significance of colors and shapes as well as the direction of the arrows in Figure  1, to facilitate clearer understanding.
In Figure  1, the various colored squares represent specific elements: blue squares denote original instructions, orange squares indicate evolved instructions, cyan squares signify model-generated solution processes, and grey squares correspond to a series of training-related operations such as supervised fine-tuning (SFT), reward modeling, and reinforcement learning (RL). To enhance the mathematical reasoning capabilities of large language models, we propose the RLEIF method, which integrates instruction evolution with reinforcement learning. This method consists of three primary steps:
  1. 1.
    Instruction Evolution and SFT
    In the first step, we apply upward and downward instruction evolution on the GSM8k and MATH datasets, generating evolved instructions for the SFT. On the leftmost side of Figure  1, the three blue arrows, from top to bottom, represent:
    1. (a)
      the adoption of the instruction evolution technique,
    2. (b)
      the generation of evolved instruction data, and
    3. (c)
      its application to SFT training.
  2. 2.
    Reward Model Training
    The second step involves two reward models: the Instruction Quality Scoring Reward Model (IRM) and the Process-Supervised Reward Model (PRM), depicted in the central section of Figure  1.
    • IRM: We employ upward and downward evolution on a seed instruction, yielding five instructions (original + evolved). These instructions are ranked by quality (e.g., C $>$ A = E $>$ B $>$ D) using GPT-4. Based on the rankings, we train the Instruction Ranking Model (IRM) to assess instruction quality. In Figure  1, this process is shown in the left-central segment: “A” represents the original instruction, while “B,” “C,” “D,” and “E” denote the evolved instructions. The first blue arrow illustrates the ranking process via GPT-4, the second arrow shows the ranking outcomes, and the third arrow highlights the use of this ranked data to train the IRM.
    • PRM: In the middle-right section of Figure  1, the process for training the PRM is depicted. The SFT model generates step-by-step solutions from the given instructions, which are then evaluated and labeled by GPT-4. This labeled data is subsequently used to train the PRM.
  3. 3.
    Reinforcement Learning with PPO
    In the final step, we integrate the IRM and PRM within a Proximal Policy Optimization (PPO)-based reinforcement learning framework. As depicted in the far-right section of Figure  1, the process is as follows:
    1. (a)
      The first blue arrow represents instruction scoring by the IRM.
    2. (b)
      The second blue arrow shows PPO initialization and the start of reinforcement.
    3. (c)
      The third blue arrow illustrates the policy model generating responses based on instructions.
    4. (d)
      The fourth blue arrow shows the scoring of each response step using the PRM.
    5. (e)
      Arrows five through eight depict the combination of IRM and PRM scores to calculate the final reward score.
    6. (f)
      The ninth blue arrow highlights the use of the reward score for the PPO training.
By integrating instruction evolution and reward-based optimization, the RLEIF method significantly enhances the reasoning capabilities of large language models.
A.7 Compare our WizardMath-SFT model across various base models (0.1B-70B) with the SOTA models on the GSM8k and Math benchmarks.
To provide a more comprehensive and fair comparison, we have included the WizardMath-SFT results in Table  13 and Table  14. These results evaluate the performance of WizardMath-SFT, trained exclusively using SFT, against current SOTA models across various base models. The key findings are summarized as follows:
  1. 1.
    Performance Comparison:
    • On Llama-2-7B and Mistral-7B-v0.1, WizardMath-SFT performs marginally below SOTA models (i.e.,Xwin-Math and Skywork-Math) and outperforms existing other excellent models (i.e.,DART-Math).
    • On Llama-2-13B and Llama-2-70B, WizardMath-SFT achieves comparable performance to Xwin-Math.
    • On all various base models, WizardMath-SFT surpasses most existing SOTA models trained solely with SFT(i.e.,DART-Math).
    Notably, WizardMath-SFT achieves these results using only 418K synthetic data points, a significantly smaller dataset compared to DART-Math (580k-590k), Xwin-Math (1440K) and Skywork-Math (2500K).
  2. 2.
    Comparison with advanced data synthesis methods (i.e., DART-Math, MetaMath)
    As shown in the following Table  15, DART-Math demonstrates strong performance across various base models and the data synthesis method proposed by DART-Math shows the effectiveness and outstanding performance. Meanwhile, WizardMath-SFT demonstrates comparable or superior performance to advanced data synthesis methods, such as DART-Math and MetaMath, across all base models. Key observations include:
    • On Mistral-7B-v0.1 and DeepSeekMath, WizardMath-SFT performs on par with DART-Math (Uniform & Prop2Diff) on GSM8k and surpasses DART-Math (Uniform & Prop2Diff) on MATH;
    • On Llama3.2 1B, Llama3.2 3B, Llama3-8B, and Llama3.1-8B, Llama2-7B, WizardMath-SFT exhibits a 2%–7% improvement over DART-Math (Uniform & Prop2Diff) on the GSM8k benchmark. On the MATH benchmark, WizardMath-SFT outperforms DART-Math (Uniform & Prop2Diff) by approximately 5% – 10%.
    These findings highlight the effectiveness of the proposed Math Evol-Instruct for enhancing mathematical reasoning capabilities.
Notably, to compare the advanced data synthesis methods such as DART-Math and MetaMath on different base models, and ensure the same training settings as in our paper during the SFT stage, we employ a learning rate of 2e-5 for the Llama series base models (i.e., Llama2 7B, Llama3.1 8B, Llama3.2 1B, and Llama3.2 3B) and a learning rate of 5e-6 for Mistral-7B-v0.1. All models are trained for 3 epochs with a batch size of 256, and 4 checkpoints are saved per epoch. Finally, we select the checkpoint with the highest accuracy on the GSM8k and MATH benchmarks for reporting.
@@
KPMath-Plus (Huang et al., 2024) Mistral-v0.1 7B 82.1 46.8
@@
KPMath-Plus (Huang et al., 2024) DeepSeekMath 7B 83.9 48.8
@@
KPMath-Plus (Huang et al., 2024) Llama 2 13B 81.6 41.0
@@
KPMath-Plus (Huang et al., 2024) Llama-2 70B 87.4 48.6
@@
Table 16: The performance of WizardMath on the GSM8k and MATH based on the Mathstral-7B-v0.1-Base, Qwen2.5-7B-Base, Qwen2.5-Math-1.5B-Base, and Qwen2.5-Math-7B-Base
Models Base Params GSM8k MATH
Mathstral-v0.1-Base - 7B 77.1 56.6
WizardMath-Mathstral Mathstral-v0.1-Base 7B 93.8 70.9
Qwen2.5-Math-Base - 1.5B 76.8 49.8
WizardMath-Qwen2.5-Math Qwen2.5-Math-Base 1.5B 86.7 68.6
Qwen2.5-Math-Base - 7B 91.6 55.4
WizardMath-Qwen2.5-Math Qwen2.5-Math-Base 7B 93.9 77.8
Qwen2.5-Base - 7B 85.4 49.8
WizardMath-Qwen2.5 Qwen2.5-Base 7B 94.0 74.5
@@
A.8 The performance of WizardMath on the other different base models
Table  16 supplements the performance improvements of Mathstral-7B-v0.1-Base, Qwen2.5-7B-Base, Qwen2.5-Math-1.5B-Base, and Qwen2.5-Math-7B-Base on the GSM8k and MATH datasets.
The results demonstrate that using Mathstral-7B-v0.1-Base as the base model, WizardMath-Mathstral improves performance by 16.7% on GSM8k (93.8 vs. 77.1) and 14.5% on MATH (70.9 vs. 56.6). When employing Qwen2.5-Math-1.5B-Base as the base model, WizardMath-Qwen2.5-Math-1.5B achieves 9.9% improvement on GSM8k (86.7 vs. 76.8) and 18.8% on MATH (68.6 vs. 49.8). Similarly, with Qwen2.5-Math-7B-Base, WizardMath-Qwen2.5-Math-7B shows a 2.3% increase on GSM8k (93.9 vs. 91.6) and 22.4% on MATH (77.8 vs. 55.4). Finally, using Qwen2.5-7B-Base as the base model, WizardMath-Qwen2.5-7B improves by 8.6% on GSM8k (94.0 vs. 85.4) and 24.7% on MATH (74.5 vs. 49.8).
Notably, both Mathstral-7B-v0.1-Base and Qwen2.5-Math-Base, pre-trained on extensive mathematical corpora, exhibit robust mathematical reasoning capabilities and deliver strong performance on GSM8k and MATH datasets. However, our proposed RLEIF method achieves substantial performance enhancements even with these highly math-optimized models. Specifically, on the MATH, RLEIF delivers a performance boost of 15% 25%, while on GSM8k, the improvement ranges from 8% 16% (with the exception of Qwen2.5-Math-7B-Base, which achieves a high baseline of 91.6 on GSM8k but still benefits from a 2.3% enhancement). These results underscore the continuous improvement enabled by our RLEIF method on models pre-trained with specialized mathematical corpora, further validating its effectiveness and scalability.